The total number of atomic orbitals in fourth energy level of an atom is
(1) 4
(2) 8
(3) 16
(4) 32

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol: Number of atomic orbitals in an orbit = n^2 = 4^2 = 16 Answer : (3)